# Are Large Language Models Capable of Generating Human-Level Narratives?

### v1

@@ Abstract
Yufei Tian1thanks:   Co-First. Our codes and the expert annotated dataset will be released at https://github.com/PlusLabNLP/Narrative-Discourse  Tenghao Huang2∗  Miri Liu1  Derek Jiang1  Alexander Spangher2 Muhao Chen23Jonathan May2Nanyun Peng1 1University of California, Los Angeles, 2University of Southern California 3University of California, Davis https://github.com/PlusLabNLP/Narrative-Discourse yufeit@cs.ucla.edutenghaoh@usc.edu
@@ 1 Introduction
Storytelling serves as an integral part in shaping our understandings of ourselves, our society and our world Langer (1942); Kaniss (1991). As large language models (LLMs) grow in capabilities Minaee et al. (2024) and are integrated into our daily communicative routines Kasneci et al. (2023), assessing the kinds of narratives they tell is crucial to understanding the ways they will shape our society.
@@
However, a computational framework remains to be established for quantitative assessments of narratives at the global or discourse level. We take a step towards the much-desired analytical framework and attempt to answer pertinent questions such as: do stories generated by LLMs exhibit the same narrative complexity and diversity as human storytelling? Do LLMs have the capacity to comprehend narrative structures? Concretely, we measure narrative discourse structures at three distinct levels: 1) story arcs (i.e., macro-level narrative development), 2) turning points (i.e., meso-level shifts) and 3) arousal & valence (i.e., micro-level dynamics). We collect a dataset of movie synopses, on which we conduct a wide range of human and automated annotations for each of these levels, with the goals of (1) contrasting LLM and human-level storytelling, and (2) probing LLM narrative structure comprehension.
@@
    We unify three levels of discourse in narrative analysis: story arc, turning point, and affective dimension. Based on this, we present the first quantitative analysis framework for narrative development, and demonstrate that it can be operationalized by humans on benchmark dataset which we release (§2 and §3).
@@
    We demonstrate that a discourse-aware generation process with LLMs (§6) — i.e. incorporating and reasoning about the story arc or turning points—enhances their overall narrative construction, as is reflected in improved suspense, emotion provocation, and narrative diversity.
@@ 3 Data Collection and Annotation
Films are our culture’s Gesamtkunstwerk (or “total work of art”) according to Michelson (1991), and our mass market vehicle for telling narratives Balio (2013). As such, we take films as a basis for exploring the stories our culture tells itself. In this section, we will describe first how we built our dataset of film plots, and then we will describe our annotation approach to study the plots’ discourse structures.
@@ 3.1 Data Preparation
We crawl the recent English-language films category on Wikipedia [^1] https://en.m.wikipedia.org/wiki/Category:2020s_English-language_films to collect the titles, genres, release dates and synopses of these films. To minimize the risk of model memorization, we filter out well-known movies using the lengths of Wikipedia pages as an approximate indicator of popularity, resulting in 819 synopses. To further avoid data contamination, we rephrase the titles and initial settings by altering all the unique identifiers such as proper nouns. Finally, we instruct GPT-4 using the rephrased titles, initial settings, and the genres to generate a paired synopsis for each collected film, resulting in 1638 synopses. All human and machine narratives are roughly of the same length.
@@ 3.2 Analysis Approaches
We seek to collect human annotations for each synopsis. To do so, we design annotation tasks for input narratives to label each with a story arc, and locate the sentential position where each of the five turning points occurs.
We recruited 16 annotators who either hold (or are pursuing) a bachelor’s degree in English or have prior experience in story analysis. To ensure the reliability of our annotators, we administered a qualification task, exemplified in Figure 10 in the appendix. We also designed short questions to assist annotators in determining the story arcs more accurately. Two example pairs of human and GPT-4 written narratives, along with corresponding human annotations, are compiled in Appendix B.1. We also detail our annotation guidelines in Appendix C.1.
We had a total of 440 narratives annotated, with each narrative annotated by three workers in-depth. The inter-annotator agreements (IAA) for the two tasks are measured at 0.90 (using Spearman’s Correlation) and 0.62 (using Cohen’s Kappa), which indicate a substantial agreement. Considering extensive labor for an in-depth human study at scale, our annotators limited their evaluations to stories produced by humans and GPT-4-0613, one of the most powerful LLMs, which should approximate the upper bound of current LLM capabilities.
@@ 5.1 Benchmark Tasks
We formulate two settings for each task: (1) we seek to identify turning points or story arcs given just the text of the narrative (2) we give the model additional discourse-level features to aid in each task. Prior research has found that additional discourse information can improve narrative reasoning Spangher et al. (2021, 2024): we hypothesize that macro-level story discourse and meso-level information are related. More specifically, for turning point identification, information about the overarching story arc type is provided. Conversely, when identifying story arcs, descriptions of key turning points within the narrative are included. To assess how well models are able to identify story arc and turning points, we compare the model’s classifications with ground truth annotations provided by human experts.
@@ 5.2 Models
We collect classifications from multiple state-of-the-art language models, including GPT-3.5, GPT-4,  [^2] April 29th. 2024 version Gemini 1.0 Pro, Claude3, Llama3-8B Touvron et al. (2023). For turning point identification, we instruct a model to analyze and tag key turning points in a movie synopsis with explanations. To enhance the model’s counting ability, all narratives are tagged with a sentence index. The exact prompts used are shown in Appendix D. We use accuracy as the metric to measure how well the models’ predictions align with expert annotations.
@@ 6 Towards Better Machine Storytelling
Next, we compare the narrative diversity between sets of Arc Enhanced stories and Outline-Only stories [^3] We instruct annotators to examine diversity in the following aspects: 1) Thematic: the central ideas conveyed; 2) Setting: when and where these stories take place; 3) Conflict type: including but not limited to character with self, other characters, society, nature, technology, fate; 4) Character: including but not limited to personality, background, development, and relationships. As seen in Table 6, Arc Enhanced significantly outperforms Outline-Only across most aspects of diversity that we considered: thematic, conflict-type, and character. We conclude that story arc discourse is a significant driver of many aspects of narrative diversity, affirming the basic truth of Vonnegut (1995)’s assertion that stories can be broadly categorized into story arc types. [^4] After all annotation tasks have been completed, we conduct two follow-up interviews with our annotators to gather their comments as qualitative feedback, which we report in Appendix B.2.
@@ 7 Related Work
Discourse-Aware Evaluation. In contrast to conventional story evaluation frameworks, which primarily focus on fluency and coherence, discourse-aware evaluation focuses on critiquing the structural and creative quality of machine-generated content. Liu et al. (2024) introduced a model that assesses stories by embedding conventional narrative structures within the evaluation process. Complementing this, Chakrabarty et al. (2024) explore narrative differences between humans and AI through a qualitative study and Li et al. (2024) reveal that RLHF-aligned language models are less diverse than the base LMs.. Begus (2023) delves into the creative outputs of LLMs, questioning their true creativity versus their capacity to merely replicate observed patterns, which encourages further exploration of the creative limits of these models. Additionally, Wang et al. (2023) introduces the Positional Discourse Coherence metric to quantitatively assess logical narrative progression.
However, prior works have been limited by the vagueness by which they have defined creativity and discourse structure. We take inspiration, in this work, from a distinct but related direction of discourse analysis: non-fictional story-telling. Van Dijk (1988) developed linguistic theories of news writing, identifying different structural elements in news articles (e.g. “Background”, “Main Event”, “Analysis”). Choubey et al. (2020); Spangher et al. (2021) showed these elements could be classified by language models and Spangher et al. (2022) used this as a lens to measure the structural output of language model generations (showing that they were dissimilar to human generations). In a similar vein, we define specific story arcs and turning points in creative stories, based on literary theory Vonnegut (1995). Likewise, we study how stories generated by LMs are structurally dissimilar to human-generated stories.
@@ 8 Conclusion
This work aims to advance the understanding and generation of narratives through the lens of three discourse elements: story arcs at macro-level, turning points at meso-level, and affective dimensions at micro-level. We contribute an expert-annotated dataset, based on which we conducted quantitative comparison between human and AI in terms of narrative generation and comprehension: LLMs fall short especially in story writing. We also show promising results that discourse-aware generation improves AI’s story-telling ability in terms of suspense, emotion engagement, and narrative diversity. We hope the collected dataset and experimental results, along with our proposed perspective, will attract wider academic interest in discourse studies and provide insights into better narrative generation, comprehension, and evaluation.
@@
Another limitation is that our research primarily focuses on English-based LLMs and resources. Future research can look into expanding this scope to include multilingual language models and diverse linguistic resources. This expansion could help to better understand and predict flavors across different cultural and linguistic contexts, potentially uncovering unique insights and flavor combinations that are specific to various cuisines and regional preferences.
@@ Appendix B Examples of Generated Narratives and Annotations
Figure 8: Human annotators’ detailed feedback on machine-generated stories when comparing the Outline-Only vs Arc-Enhanced strategy. They are blind to the prompting strategies and all presented stories were randomly shuffled. We reconstruct their comments and color-code with green for favorable ones and red for unfavorable ones.
Figure 9: Figure 8 Continued.

### v2

@@ Abstract
Yufei Tian1thanks:   The two authors contributed equally.  Tenghao Huang2∗  Miri Liu1  Derek Jiang1  Alexander Spangher2 Muhao Chen23Jonathan May2Nanyun Peng1 1University of California, Los Angeles, 2University of Southern California 3University of California, Davis https://github.com/PlusLabNLP/Narrative-Discourse yufeit@cs.ucla.edutenghaoh@usc.edu
@@ 1 Introduction
Storytelling serves as an integral part in shaping our understandings of ourselves, our society and our world Langer (1942); Kaniss (1991). As large language models (LLMs) grow in capabilities Minaee et al. (2024) and are integrated into our daily communicative routines Kasneci et al. (2023), assessing the narrative structures of the stories they tell is crucial to understanding the ways they will shape our society.
@@
However, a computational framework remains to be established for quantitative assessments of narratives at the global or discourse level. We take a step towards the much-desired analytical framework and attempt to answer pertinent questions such as: do stories generated by LLMs exhibit the same narrative complexity and diversity as human storytelling? Do LLMs have the capacity to comprehend narrative structures? Concretely, we measure narrative discourse structures at three distinct levels: 1) story arcs (i.e., macro-level narrative development), 2) turning points (i.e., meso-level shifts) and 3) arousal & valence (i.e., micro-level dynamics). We collect a dataset of movie synopses, on which we conduct a wide range of human and automated annotations for each of these levels, with the goals of (1) contrasting LLM and human storytelling, and (2) probing LLM narrative structure comprehension.
@@
    We unify three levels of discourse in narrative analysis: story arc, turning point, and affective dimension. Based on this, we present the first quantitative analysis framework to benchmark narrative development, and demonstrate that it can be operationalized by humans on benchmark dataset which we release (§2 and §3).
@@
    We demonstrate that a discourse-aware generation process with LLMs (§6) — i.e. incorporating and reasoning about the story arc or turning points—enhances their overall narrative construction, as is reflected in improved suspense, emotion provocation, and narrative diversity. This lays the groundwork for future research to refine models to incorporate complex narrative structures for storytelling and beyond.
@@ 3 Data Collection and Annotation
Films are our culture’s Gesamtkunstwerk (or “total work of art”) according to Michelson (1991), and our mass market vehicle for telling narratives Balio (2013). Thus, we take films as a basis for exploring the stories our culture tells itself. The narratives we examine are condensed versions of some of the most intricate storylines humans create—those found in movies. While these synopses focus on key plot developments, they should not be considered simple or straightforward. In this section, we will describe first how we built our dataset of film plots, and then we will describe our annotation approach to study the plots’ discourse structures.
@@ 3.1 Data Preparation
We crawl the recent English-language films category on Wikipedia [^1] https://en.m.wikipedia.org/wiki/Category:2020s_English-language_films to collect the titles, genres, release dates and synopses of these films. To avoid data leakage, we filter out well-known movies using the lengths of Wikipedia pages as an approximate indicator of popularity, resulting in 819 synopses. To further avoid data contamination, we rephrase the titles and initial settings by altering all the unique identifiers such as proper nouns. Finally, we instruct GPT-4 using the rephrased titles, initial settings, and the genres to generate a paired synopsis for each collected film, resulting in 1638 synopses. All human and machine narratives are roughly of the same length.
@@ 3.2 Analysis Approaches
We seek to collect human annotations for each synopsis. To do so, we design annotation tasks for input narratives to label each with a story arc, and locate the sentential position where each of the five turning points occurs. We introduce a few key modifications to the turning point schema of Papalampidi et al. (2019) to handle more complex narratives: 1) flexible positioning for TP3 and TP4, and 2) optional, though discouraged, missing or multiple positions for TP2, TP3, and TP4 when annotators are uncertain.
We recruited 16 annotators who either hold (or are pursuing) a bachelor’s degree in English or have prior experience in story analysis. To ensure the reliability of our annotators, we conducted multiple training sessions to fully onboard our annotators and administered a qualification task, exemplified in Figure 10 in the appendix. We also designed short questions to assist annotators in determining the story arcs more accurately. Two example pairs of human and GPT-4 written narratives, along with corresponding human annotations, are compiled in Appendix B.1. We also detail our annotation guidelines in Appendix C.1.
The narratives we studied have an average of 705.6 words and 37.8 sentences, longer and more complex compared to most other papers quantitatively studying narratives, such as Sap et al. (2022) with 300 words and Li et al. (2018) with 18.5 sentences. We had a total of 440 narratives annotated, with each narrative annotated by three workers in-depth. The inter-annotator agreements (IAA) for the two tasks are measured at 0.90 (using Spearman’s Correlation) and 0.62 (using Cohen’s Kappa), which indicate a substantial agreement and speaks to the quality of our annotation process. Considering extensive labor for an in-depth human study at scale, our annotators limited their evaluations to stories produced by humans and GPT-4-0613, one of the most powerful LLMs, which should approximate the upper bound of current LLM capabilities.
@@ 5.1 Benchmark Tasks
We formulate two settings for each task: (1) we seek to identify turning points or story arcs given just the text of the narrative (2) we give the model additional discourse-level features to aid in each task. Prior research has found that additional discourse information can improve narrative reasoning Spangher et al. (2021, 2024a): we hypothesize that macro-level story discourse and meso-level information are related. More specifically, for turning point identification, information about the overarching story arc type is provided. Conversely, when identifying story arcs, descriptions of key turning points within the narrative are included. To assess how well models are able to identify story arc and turning points, we compare the model’s classifications with ground truth annotations provided by human experts.
@@ 5.2 Models
We collect classifications from multiple state-of-the-art language models, including GPT-3.5, GPT-4,  [^2] April 29th. 2024 version Gemini 1.0 Pro, Claude3, Llama3-8B Touvron et al. (2023). For turning point identification, we instruct a model to analyze and tag key turning points in a movie synopsis with explanations. To enhance the model’s counting ability, all narratives are tagged with a sentence index. The exact prompts used are shown in Appendix D.2. We use accuracy as the metric to measure how well the models’ predictions align with expert annotations.
@@ 6 Towards Better Machine Storytelling
Next, we compare the narrative diversity between sets of Arc Enhanced stories and Outline-Only stories [^3] We instruct annotators to examine diversity in the following aspects: 1) Thematic: the central ideas conveyed; 2) Setting: when and where these stories take place; 3) Conflict type: including but not limited to character with self, other characters, society, nature, technology, fate; 4) Character: including but not limited to personality, background, development, and relationships. As seen in Table 6, Arc Enhanced significantly outperforms Outline-Only across most aspects of diversity that we considered: thematic, conflict-type, and character. We conclude that story arc discourse is a significant driver of many aspects of narrative diversity, affirming the basic truth of Vonnegut (1995)’s assertion that stories can be broadly categorized into story arc types.
7 Qualitative Study: Understanding Human Preferences
After completing all annotation tasks, we conducted two follow-up interviews to gather qualitative feedback from our annotators. These interviews focused on annotators who worked on the task in §$\S$ 6 — reading pairs of machine generated narratives (Outline-Only vs Arc-Enhanced) and examining diversities plus overall preference. They were encouraged to freely provide justifications or comments on any of the readings. After the interviews, we reconstructed their feedback and presented representative comments in Figure 7 and Figure 9. Overall, the human annotators prefer concrete narratives with twists in plot development that are logical and well-motivated. They dislike straightforward, positive plots or those with ‘miracle turns’ that are not adequately justified.
Figure 7: Human annotators’ feedback on machine-generated stories when comparing the Outline-Only vs Arc-Enhanced strategy. They were blind to the prompting strategies and all presented stories were randomly shuffled. We reconstruct their comments and color-code with green for favorable ones and red for unfavorable ones. Continued in Figure 9.
@@ 8 Related Work
Discourse-Aware Evaluation. In contrast to conventional story evaluation frameworks, which primarily focus on fluency and coherence, discourse-aware evaluation focuses on critiquing the structural and creative quality of machine-generated content Harel-Canada et al. (2024); Spangher et al. (2024b); Tian et al. (2024). Liu et al. (2024) introduced a model that assesses stories by embedding conventional narrative structures within the evaluation process. Complementing this, Chakrabarty et al. (2024) explore narrative differences between humans and AI through a qualitative study and Li et al. (2024) reveal that RLHF-aligned language models are less diverse than the base LMs.. Begus (2023) delves into the creative outputs of LLMs, questioning their true creativity versus their capacity to merely replicate observed patterns, which encourages further exploration of the creative limits of these models. Additionally, Wang et al. (2023) introduces the Positional Discourse Coherence metric to quantitatively assess logical narrative progression.
However, prior works have been limited by the vague definitions of creativity and discourse structure. We take inspiration from the literature theories of discourse analysis of drama and fiction Labov and Waletzky (1967); Vonnegut (1995). For instance, Freytag (1894) proposed a a five-part dramatic structure, now commonly understood as Exposition, Rising Action, Climax, Falling Action, and Dénouement. Li et al. (2018) combined such literature theories and annotated story macro-structures. In non-fiction storytelling like news writing, Choubey et al. (2020); Spangher et al. (2021, 2022) demonstrated that language models can classify similar elements, with artificial news differing from human-generated content. Likewise, we define specific story arcs and turning points in creative stories and examine how stories generated by LMs structurally differ from human ones.
@@ 9 Conclusion
This work aims to advance the understanding and generation of narratives through the lens of three discourse elements: story arcs at macro-level, turning points at meso-level, and affective dimensions at micro-level. We contribute an expert-annotated dataset, based on which we conducted quantitative comparison between human and AI in terms of narrative generation and comprehension: LLMs fall short especially in story writing. We find models lack narrative diversity, and struggle at develop crucial turning points, such as major setback and climax, leading to less engaging stories. We also show promising results that discourse-aware generation improves AI’s story-telling ability in terms of suspense, emotion engagement, and narrative diversity.
We view our effort as a useful starting point towards a systematic analysis of narrative discourse. We hope the collected dataset and experimental results, along with our proposed perspective, will attract wider academic interest in discourse studies and provide insights into better narrative generation, comprehension, and evaluation.
@@
Another limitation is that our research primarily focuses on English-based LLMs and resources. Our initial focus on English allows us to establish the discourse-level framework before expanding to others. Future research can look into expanding this scope to include multilingual language models and diverse linguistic resources. This expansion could help to better understand and predict flavors across different cultural and linguistic contexts, potentially uncovering unique insights and flavor combinations that are specific to various cuisines and regional preferences.
Earlier studies, such as Huang et al. (2021), have identified the presence of gender biases within this dataset. Consequently, we want to point out that generating stories using our pipeline may also be at risk of perpetuating and intensifying these biases.
Acknowledgement
The authors would like to thank the PlusLab members at UCLA and the anonymous reviewers for their valuable feedback and helpful discussions. This research is partly supported by National Science Foundation CAREER award #2339766, an Amazon AGI foundation research award, and a Google Research Scholar grant. Tenghao Huang and Jonathan May are supported by the Defense Advanced Research Projects Agency with award HR00112220046.
@@ Appendix B Examples of Generated Narratives and Annotations
Figure 9: Figure 7 Continued. Prior to the interviews and after the all annotation tasks were completed, annotators were already informed that the presented stories were generated by AI, but the specific methods and models were hidden from them.
@@ Appendix D Prompt details
D.1 Prompts Used in Data Preparation
We consider the introductory part (first 1-3 sentences in the human-written narrative) as the initial setting. We asked GPT-4 to rephrase the setting by replacing all proper nouns (names, places, anything unique), and then change the phrasing slightly to return the new setting. Based on the newly rephrased initial setting and the original title, we asked the model to generate a similar but not identical title.
D.2 Prompts Used in Benchmarking